A contractor hires two trucking companies to haul 1600 tons of crushed stone for a highway construction project. The contracts state that one company is to haul four times as much as the other. Find the amount hauled by each company.

Answer by checkley77(12844)

4X+X=1600
5X=1600
X=1600/5
X=320 TONS IS THE SMALL LOAD.
1600-320=1280 IS THE LARGER LOAD.
PROOF:
4*320+320=1600
1280+320=1600